# Black Hole Search by a Set of Scattered Agents in Dynamic Rings

### v1

@@ Abstract
In this paper we investigate the problem of searching for a black hole in a dynamic graph by a set of scattered agents (i.e., the agents start from arbitrary locations of the graph). The black hole is a node that silently destroys any agent visiting it. This kind of malicious node nicely models network failures such as a crashed host or a virus that erases the visiting agents. The black hole search problem is solved when at least one agent survives, and it has the entire map of the graph with the location of the black hole. We consider the case in which the underlining graph is a dynamic 1-interval connected ring: a ring graph in which at each round at most one edge can be missing. We first show that the problem cannot be solved if the agents can only communicate by using a face-to-face mechanism: this holds for any set of agents of constant size, with respect to the size $n$ of the ring.
To circumvent this impossibility we consider agents equipped with movable pebbles that can be left on nodes as a form of communication with other agents. When pebbles are available, three agents can localize the black hole in $O(n^{2})$ moves. We show that such a number of agents is optimal. We also show that the complexity is tight, that is $\Omega(n^{2})$ moves are required for any algorithm solving the problem with three agents, even with stronger communication mechanisms (e.g., a whiteboard on each node on which agents can write messages of unlimited size). To the best of our knowledge this is the first paper examining the problem of searching a black hole in a dynamic environment with scattered agents.
@@ 1.1 Exploration of Dynamic Networks
In the distributed computing community a large set of works (see [20]) studies the computational paradigm of mobile agents. A mobile agent is a software component that is able to move on a network visiting nodes. When a node is visited, the agent executes some computation interacting with the local environment, the memory and resources of the visited computational node, and then it moves on a next computational node by transmitting itself on the network. That is the agent can be seen as an intelligent message with computational capabilities and that is able to decide its next destination.
@@
In the mobile agent paradigm a plethora of problems have been studied. The most famous are: exploration, the agents have to collectively visit the entire network; gathering, the agents have to reach the same node; patrolling, the agents have to periodically patrol the network minimising the time between two visits of the same node.
One problem that has been thoroughly investigated is the Black Hole Search (Bhs) [25]. In this problem there exists a dangerous stationary node, called black hole (Bh), that silently erases from the network all the agents that visit it. A Bh node could model several kind of common failures: for instance, consider a crashed host, and any agent trying to visit it will be lost and removed from the network; or a node infected by a virus that cancels the incoming agents. The many works that investigated Bh have given to us a complete knowledge of the computational properties of the problem under several assumptions (examples are communication mechanisms employed by the agents, level of synchronicity, topology of the network, and agents’ knowledge and capabilities). However, almost all of them examine the case in which the network is static: the set of computational nodes and the links connecting them are static and never change.
@@
Recently, research within distributed computing has started to focus on mobile-agents in highly dynamic graphs, i.e., graphs where the topological changes are not limited to sporadic and disruptive events (such as process failures, links congestion, etc). Highly dynamic graphs model a wide range of modern networked systems whose dynamic nature is the natural product of innovations in communication technology (e.g., wireless networks), in software layer (e.g., a controller in a software defined network), and in society (e.g., the pervasive nature of smart mobile devices) (e.g., see [5]).
@@
We consider evolving graphs, that are dynamic graphs that can be seen as an infinite sequence of static graphs. In this case the model of computation is by definition synchronous, and at each round corresponds a static graph of the aforementioned sequence. A popular assumption in this model is the 1-interval connectivity: this assumption dictates that at each round the dynamic graph is connected (e.g., [1, 10, 21, 23, 22, 26]).
@@
We focus on the case of 1-interval connected rings, where the topology is a ring graph in which at each round at most one edge is missing.
In the last years the body of works studying agents on highly dynamic graphs has been growing with a sustained velocity (for a recent survey see [9]). In particular, a large number of papers is focussing on 1-interval connected rings: the gathering problem has been investigated in [12], the exploration problem in [11, 3, 4] and the Bhs for colocated [^1] Agents are colocated when they all start from the same node. agents in the recent [13]. Despite this large interest, a lot of questions are open. One of them is answered in this paper: how does the computational landscape of finding a Bh change when agents are scattered? In the scattered case agents start from arbitrary nodes of the graph. We will show that this setting has many differences with respect to the case of colocated agents (studied in [13]) in terms of both solvability (solving Bhs in some settings becomes impossible) and complexity.
@@ 1.2 Related Works
[13] studies the Bhs in the same setting studied in this paper: 1-interval connected ring with a single Bh. [13] shows that three agents are necessary to find the Bh (in the static case two agents are enough to explore arbitrary known graphs [7]), and it presents optimal algorithms to find the Bh with three agents. Moreover, if agents can communicate only when they are on the same node, the authors show that Bhs can be solved in ${\Theta}(n^{2})$ moves and rounds. Finally, if agents can use pebbles, they show an improved algorithm that finds the Bh in ${\Theta}(n^{1.5})$ moves and rounds.
@@ 1.3 Contributions
We study the problem of finding a Bh in an oriented dynamic ring by a set of scattered agents with visible identities under different communication capabilities. We study two main families of communication mechanisms. The endogenous family, where the agents can communicate without using any external facilities. In this case, they can either see each other only when on the same node (Vision model), or they can also talk with each other (Face2Face model). In contrast, in the exogenous family, the agents communicate using external tools. In this case, we have the Pebbles model, in which each agent carries a pebble that can leave on a node to mark it for other agents, or that it can remove from a node in case this marking is not needed anymore; or the Whiteboard model, in which each node has a public whiteboard on which agents can write messages of unlimited size.
Our first result (Obs. 2) is that in the endogenous family the Bhs is unsolvable using three agents. This is in contrast with the colocated case where Bhs is solvable in the same setting by using three agents (see [13]).
To circumvent such impossibility we then rely on exogenous communication. We focus on algorithms that use an optimal number of agents. In particular, in [13], it has been shown that Bhs in dynamic rings is unsolvable if only two agents are available. This results clearly extends also to the scattered case. Therefore, we will consider algorithms for set of three agents (size optimal algorithms). In Th. 4, we show that any optimal size algorithm solving Bhs requires $\Omega(n^{2})$ moves and $\Omega(n^{2})$ rounds in the whiteboard model. We note that on a static and synchronous rings two agents can find the blackhole in $O(n)$ moves and rounds. This observation is interesting as it shows that dynamicity does not only increase the number of required agents but it also increases, significatively, the time and moves required. This also highlights the price to pay for having scattered agents: in fact, for dynamic rings, $O(n^{1.5})$ moves and rounds are sufficient in the colocated case [13]. Finally, our lower bound is tight: we provide an algorithm that solves Bhs in the pebble model in $O(n^{2})$ rounds and moves using three agents (Th. 6).
@@ 2.1 The Model and the Problem
We consider two classes of communication mechanisms (endogenous and exogenous) which give rise to four models.
Endogenous Mechanisms rely only on the robots’ capabilities without requiring any external object. In the FaceToFace (F2F) model the agent can explicitly communicate among themselves only when they reside on the same node. In the Vision model an agents can see all the other agents that reside on the same node (hence count their number); however, they cannot communicate.
Exogenous Mechanisms do require external objects for the robots to exchange information. Among those we distinguish:
@@ 3.4 CautiousPendulum: An algorithm for colocated agents
4 Impossibility with Scattered Agents and Endogenous Communication
When the agents are scattered, three of them, even equipped with the stronger endogenous mechanism (i.e., F2F model), cannot solve Bhs on rings of arbitrary size, as shown by the following:
Observation 2.
Proof.
Fortunately, any exogenous mechanism circumvents the impossibility of Obs. 2. In the following we focus on such mechanisms.
5 Exogenous Communications: Lower Bound for Size-optimal Algorithms.
@@ 4 Impossibility with Scattered Agents and Endogenous Communication
6 An Optimal Exogenous Algorithm: Gather&Locate
@@
6.2 Correctness of Gather&Locate.
@@
  • One agent terminates, and by Observation 3 it terminates correctly solving the BHS-Problem. The other agent has to still locate the Bh
@@
7 Conclusion
In this paper, we have investigated the Bhs on dynamic rings in scenarios where agents are scattered. However, several questions remain open. In particular, we believe that the findings in Observation 2 could potentially be applied to any constant number of agents. Another unresolved issue is the examination of unoriented rings with scattered and anonymous agents, as well as the development of faster Bhs algorithm for groups larger than three agents.

### v3

# Black Hole Search by Scattered Agents in Dynamic Rings

@@ Abstract
In this paper, we address the challenge of locating a black hole within a dynamic graph using a team of mobile agents that start from arbitrary positions in the graph. A black hole is a node that silently eliminates any agent that visits it, effectively modeling network failures such as a crashed host or a destructive local virus. The black hole search problem is considered solved when at least one agent survives and possesses a complete map of the graph, including the precise location of the black hole.
Our study focuses on the scenario where the underlying graph is a dynamic 1-interval connected ring: a ring graph where, in each round, one edge may be absent. Agents communicate with other agents using movable pebbles that can be placed on nodes. In this setting, we demonstrate that three agents are sufficient to identify the black hole and this can be done in $O(n^{2})$ moves. Furthermore, we prove that this number of agents is optimal. Additionally, we establish that the complexity bound is tight: any algorithm solving the problem with three agents requires $\Omega(n^{2})$ moves, even when stronger communication mechanisms, such as unlimited-size whiteboards on nodes, are available.
@@ 1.1 Exploration of Dynamic Networks
In the field of distributed computing, a substantial body of research (see [20]) has focused on the computational paradigm of mobile agents. A mobile agent is a software entity capable of navigating a network by visiting nodes. Upon visiting a node, the agent executes computations while interacting with the local environment, including the memory and resources of the node, before transmitting itself to a neighboring node. Essentially, a mobile agent can be regarded as an intelligent message with computational abilities, capable of determining its next destination.
@@
Numerous problems have been studied under the mobile agent paradigm. Notable examples include: exploration, where a team of agents must collectively traverse the entire network; gathering, where agents converge at a common, initially unknown node; and patrolling, where agents periodically monitor the network, minimizing the interval between consecutive visits to the same node.
One well-studied problem in this domain is the Black Hole Search (Bhs) [25]. This involves identifying a dangerous stationary node, termed the black hole (Bh), which silently destroys any agent that visits it. A Bh models various network failures, such as a crashed host that eliminates visiting agents or a virus-infected node that deletes incoming messages. Extensive research on the Bh problem has provided a comprehensive understanding of its computational properties under different assumptions, including communication mechanisms, synchronicity levels, network topology, and agent capabilities. However, most studies have considered networks to be static, where nodes and links remain unchanged.
@@
Recently, distributed computing research has shifted towards studying mobile agents in highly dynamic graphs, where topological changes occur frequently and are not limited to sporadic failures or congestion. Such graphs represent various modern networks whose dynamic nature arises from advancements in communication technologies (e.g., wireless networks), software architectures (e.g., software-defined network controllers), and societal changes (e.g., ubiquitous smart mobile devices) (see [5]).
@@
We investigate evolving graphs, a class of dynamic graphs represented as infinite sequences of static graphs. In this model, computation is inherently synchronous, with each round corresponding to a static graph in the sequence. A widely adopted assumption in this setting is 1-interval connectivity, which ensures that the graph remains connected at every round (e.g., [1, 10, 21, 23, 22]).
@@
Our focus is on 1-interval connected rings, where the network topology is a ring graph where, in each round, an edge might be missing.
In recent years, the study of agents on highly dynamic graphs has grown rapidly (see the survey in [9]). Significant progress has been made on 1-interval connected rings, with investigations into the gathering problem [12], the exploration problem [11, 3, 4], and the Bhs for colocated [^1] Agents are colocated when they all start from the same node. agents [13].
Despite this progress, many questions remain open. This paper addresses one such question: how does the computational complexity of finding a Bh change when agents are scattered ? In the scattered scenario, agents start from arbitrary nodes in the graph. We demonstrate that this setting differs significantly from the colocated case (studied in [13]) in terms of both solvability (some scenarios become unsolvable) and computational complexity.
@@ 1.2 Related Works
In [13], the authors study the Bhs in the same setting of this paper: 1-interval connected ring with a single Bh. [13] shows that three agents are necessary to find the Bh (in the static case two agents are enough to explore arbitrary known graphs [7]), and it presents optimal algorithms to find the Bh with three agents. Moreover, if agents can communicate only when they are on the same node, the authors show that Bhs can be solved in ${\Theta}(n^{2})$ moves and rounds. Finally, if agents can use pebbles, they show an improved algorithm that finds the Bh in ${\Theta}(n^{1.5})$ moves and rounds.
@@ 1.3 Contributions
We study the problem of finding a Bh in an oriented dynamic ring by a set of scattered agents with visible identities under different communication capabilities. We study the exogenous family of communication mechanisms, where the agents communicate using external tools. In this case, we have the Pebbles model, in which each agent carries a pebble that can leave on a node to mark it for other agents, or that it can remove from a node in case this marking is not needed anymore; or the Whiteboard model, in which each node has a public whiteboard on which agents can write messages of unlimited size.
We focus on algorithms that use an optimal number of agents. In particular, in [13], it has been shown that Bhs in dynamic rings is unsolvable if only two agents are available. This results clearly extends also to the scattered case. Therefore, we will consider algorithms for set of three agents (size optimal algorithms). In Th. 4, we show that any optimal size algorithm solving Bhs requires $\Omega(n^{2})$ moves and $\Omega(n^{2})$ rounds in the whiteboard model. We note that on a static and synchronous rings two agents can find the blackhole in $O(n)$ moves and rounds. This observation is interesting as it shows that dynamicity does not only increase the number of required agents but it also increases, significatively, the time and moves required. This also highlights the price to pay for having scattered agents: in fact, for dynamic rings, $O(n^{1.5})$ moves and rounds are sufficient in the colocated case [13]. Finally, our lower bound is tight: we provide an algorithm that solves Bhs in the pebble model in $O(n^{2})$ rounds and moves using three agents (Th. 6).
We conclude the paper with on observation on the impossibility of solving the problem by three agents when using an endogenous communication mechanism, such as the Face-to-Face.
@@ 2.1 The Model and the Problem
We consider the Exogenous communication mechanism, which requires external objects for the robots to exchange information. Among those we distinguish:
@@ 3.4 CautiousPendulum: An algorithm for colocated agents
4 Lower Bound for Size-optimal Algorithms
@@
5 An Optimal Algorithm: Gather&Locate
@@ 5.1 Detailed Description
5.2 Correctness of Gather&Locate
@@
  • One agent terminates, and by Observation 2 it terminates correctly solving the BHS-Problem. The other agent has to still locate the Bh
@@
6 Observation on Endogenous Communication
We conclude the paper with an observation on the communication mechanism used by the agents.
In this paper, we have considered agents which communicate using exogenous mechanisms (i.e., provided by the system), pebbles or whiteboards.
Another family of communication mechanisms is the endogenous one, which relies only on the robots’ capabilities without requiring any external objects. In the FaceToFace (F2F) model the agent can explicitly communicate among themselves only when they reside on the same node. In the Vision model an agents can see all the other agents that reside on the same node (hence count their number); however, they cannot communicate.
As we have shown, three scattered agents equipped with the weakest exogenous communication (i.e., pebbles) suffice to solve Bhs. In the case of endogenous communication, this is no longer possible: three agents using the strongest endogenous mechanism (i.e., F2F) cannot solve Bhs on rings of arbitrary size, even if the ring is static.
Observation 3.
Proof.
Further observe that this impossibility is in contrast with the colocated case where Bhs is solvable in the same setting by using three agents (see [13]).
This leaves open the question of whether the problem is solvable with more than three agents. We suspect that the findings in Observation 3 could potentially be applied to any constant number of agents.
7 Concluding Remarks
In this paper, we have investigated the Bhs on dynamic rings in scenarios where agents with exogenous communication mechanism are scattered. Specifically, we provided solutions that are both time and size optimal.
An unresolved issue is the examination of unoriented rings with scattered and anonymous agents, as well as the development of faster Bhs algorithm for groups larger than three agents.
Acknowledgments
A preliminary version of this paper has been presented at the 27th Conference on Principles of Distributed Systems (OPODIS 2023). This research has been supported in part by NSERC under the Discovery Grant program, and by the Italian National Group for Scientific Computation GNCS-INdAM.